OpenStudy (anonymous):

Approximate the instantaneous rate of change for the functions. f(x)=-2x^2+x+4 at x=4

OpenStudy (luigi0210):

Isn't this just finding slope?

OpenStudy (anonymous):

Do I find the derivative of -2x^2

OpenStudy (anonymous):

Differentiate the function -2x^2+x+4 Then place x=4 into the derivative

OpenStudy (anonymous):

(a*b)'=a'*b+a*b', and let a=-2, b=x^2